#4 only pleaseZOOM + 3. Suppose that Firm 1 and Firm 2 are Cournot competitors. Each still has a marginal cost of 2 and the inverse demand
Answer #1

Market demand is given by P = 5 - 0.5(Q1 + Q2).

Now we have the profit function for firm 1 as π1 = (P - MC)Q1 and for firm 2 as π2 = (P - MC)Q2. This gives

π1 = (5 - 0.5Q1 - 0.5Q2 - 1)Q1 and π2 = (5 - 0.5Q1 - 0.5Q2 - 2)Q2

π1 = 4Q1 - 0.5Q1^2 - 0.5Q1Q2 and π2 = 3Q2 - 0.5Q2^2 - 0.5Q1Q2

Maximize profits and keep π1'(Q1) = π2'(Q2) = 0

4 - Q1 - 0.5Q2 = 0 and 3 - Q2 - 0.5Q1 = 0

Q1 = 4 - 0.5Q2 and Q2 = 3 - 0.5*(4 - 0.5Q2)

This gives Q2 = 1 + 0.25Q2 or Q2 = 1.33 and so Q1 = 3.33

New price = 5 - 0.5*(3.33 + 1.33) = $2.67

Profit for firm 1 = (2.67 - 1)*3.33 = $5.56 and for firm 2 profit = (2.67 - 2)*1.33 = $0.88